OpenStudy (anonymous):

Find the center, vertices and foci of 16x^2-9y^2+144=0

OpenStudy (tkhunny):

Have you considered writing it in a familiar form? \(16x^{2} - 9y^{2} + 144 = 0\) \(9y^{2} - 16x^{2} = 144\) -- Okay, it look like an hyperbola with center (0,0) \(\dfrac{y^{2}}{16} - \dfrac{x^{2}}{9} = 1\) -- Okay, it has a major axis of lenglt 2*4 in the y-direction and a minor axis length of 2*3 in the x-direction. All that from just putting it in a familiar form. Now what?

OpenStudy (anonymous):

i dont know how to get the center of it

OpenStudy (tkhunny):

This is why you should be acquainted with the FORM. \(\dfrac{(y-k)^{2}}{a^{2}} - \dfrac{(x-h)^{2}}{b^{2}} = 1\) Center is (h,k) - Every time. We have zero (0) in both cases in your problem.